OpenStudy (anonymous):

HW: d^2(y)/dx^2 for x^2 - y^2 = 5

OpenStudy (luigi0210):

Start by finding the first derivative, what do you get?

OpenStudy (anonymous):

one sec...

OpenStudy (anonymous):

2x?

OpenStudy (luigi0210):

No >_>

OpenStudy (anonymous):

I came up with that because the derivative of -y^2 is 0

OpenStudy (anonymous):

he, im lost :P

OpenStudy (anonymous):

2x-2y(y')=0 y'=-2x/-2y=x/y

OpenStudy (anonymous):

it cant be x/y because its for x^2 - y^2 = 5.

OpenStudy (luigi0210):

That was just the first derivative.

OpenStudy (luigi0210):

Now for the second we know that: \[\LARGE \frac{dy}{dx}=\frac{x}{y}\]

OpenStudy (luigi0210):

So now find: \[\LARGE \frac{d^2y}{dx^2}\]

OpenStudy (anonymous):

im confused

OpenStudy (luigi0210):

On finding the derivative?

OpenStudy (anonymous):

the second part d^2y/dx^2

OpenStudy (luigi0210):

You have to use the quotient rule

OpenStudy (anonymous):

gosh my brain is spinning

OpenStudy (anonymous):

i have to finish this paper to study on

OpenStudy (luigi0210):

Okay, I'll go step-by-step and see if that helps you?

OpenStudy (anonymous):

ok

OpenStudy (luigi0210):

Okay, so we have: \[\LARGE x^2-y^2=5\] Now, we differentiate with respect to x: \[\LARGE x^2(\frac{dy}{dx})-y^2(\frac{dy}{dx})=5(\frac{dy}{dx})\] And get: \[\LARGE 2x-2y(\frac{dy}{dx})=0\] Now get the y' by itself: \[\LARGE -2y(\frac{dy}{dx})=-2x\] Divide by -2y and simplify: \[\LARGE \frac{dy}{dx}=\frac{-2x}{-2y}=\frac{x}{y}\] Make sense so far?

OpenStudy (anonymous):

kinda so what after that?

OpenStudy (luigi0210):

Now we have to find the derivative of x/y, so we have to use the quotient rule: \[\LARGE \frac{d^2y}{dx^2}=\frac{(y*x(\frac{dy}{dx}))-(y(\frac{dy}{dx})*x)}{y^2}\] Simplify: \[\LARGE \frac{d^2y}{dx^2}=\frac{(y*1)-(\frac{dy}{dx}*x)}{y^2}\] Now we know from earlier that: \[\LARGE \frac{dy}{dx}=\frac{x}{y}\] So plug that in: \[\LARGE \frac{d^2y}{dx^2}=\frac{y-(\frac{x}{y}*x)}{y^2}\] Can you finish it?

OpenStudy (anonymous):

no it looks really herd :| its gonna be like 5 somethin' mate. i can get that

OpenStudy (anonymous):

is it -5/y?

OpenStudy (luigi0210):

Did you at least try? >_>

OpenStudy (luigi0210):

@Isaiah.Feynman Can you check my work? I think I might of messed up..

OpenStudy (anonymous):

im not sure if iget it

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im not sure how to get the final answ

OpenStudy (luigi0210):

I explained most everything .-.

OpenStudy (anonymous):

ya, but im still not sure about the final answ

OpenStudy (luigi0210):

Well if you did all the work you should of gotten something like \[\LARGE y^3-x^2y\]

OpenStudy (anonymous):

i dont think

OpenStudy (luigi0210):

Actually, I'm not 100% sure..

OpenStudy (anonymous):

how do you make it all pretty like that luigi0210 (y^2)-(x^2) divided by y^3 is what I got for the second derivative

OpenStudy (anonymous):

its gotta be like 5 somethin'

OpenStudy (luigi0210):

@cc3 \[\LARGE LaTex\] my friend~

OpenStudy (anonymous):

its like 5 divided by somethin'

OpenStudy (anonymous):

love ya mate i think i got its gonna be mighty helpful for tomorrows quiz

OpenStudy (luigi0210):

-.-